### v1

# Dual Test-time Training for Out-of-distribution Recommender System

Abstract
Deep learning has been widely applied in recommender systems, which has achieved revolutionary progress recently. However, most existing learning-based methods assume that the user and item distributions remain unchanged between the training phase and the test phase. However, the distribution of user and item features can naturally shift in real-world scenarios, potentially resulting in a substantial decrease in recommendation performance. This phenomenon can be formulated as an Out-Of-Distribution (OOD) recommendation problem. To address this challenge, we propose a novel Dual Test-Time-Training framework for OOD Recommendation, termed DT3OR. In DT3OR, we incorporate a model adaptation mechanism during the test-time phase to carefully update the recommendation model, allowing the model to specially adapt to the shifting user and item features. To be specific, we propose a self-distillation task and a contrastive task to assist the model learning both the user’s invariant interest preferences and the variant user/item characteristics during the test-time phase, thus facilitating a smooth adaptation to the shifting features. Furthermore, we provide theoretical analysis to support the rationale behind our dual test-time training framework. To the best of our knowledge, this paper is the first work to address OOD recommendation via a test-time-training strategy. We conduct experiments on three datasets with various backbones. Comprehensive experimental results have demonstrated the effectiveness of DT3OR compared to other state-of-the-art baselines.
Index Terms:
Out-Of-Distribution; Recommender System; Test-Time-Training; User/item Feature Shift
I Introduction
Recommender systems play a crucial role in alleviating information overload on social media platforms by providing personalized information filtering. In recent years, a plethora of recommendation algorithms have been proposed, including collaborative filtering [1, 2, 3, 4, 5, 6], graph-based recommendation [7, 8, 9], cross-domain recommendation [10] and etc.
Although promising performance has been achieved, most existing recommendation algorithms assume that the user and item distributions remain unchanged between the training phase and the test phase, adhering to the assumption of independent and identically distributed (I.I.D.) variables. However, these distributions often undergo shifts in many real-world scenarios, violating this assumption. The shifts primarily manifest in two aspects: user feature shift and item feature shift. User feature shift specifically refers to changes in a user’s preferences over time as his or her characteristics evolve. For example, as a boy grows older, his interests may shift from comic books to sports, such as basketball. On the other hand, item feature shift pertains to changes in preference relationships resulting from the replacement or updating of specific types of items. For instance, when items are renewed or updated, they may have new characteristics that lead to different preferences and referral relationships among users. We demonstrate the user/item feature shift in Fig. 1. These shifts significant impact on recommendation performance [11, 12]. Consequently, it is crucial to consider out-of-distribution (OOD) recommendation to address this issue.
Figure 1: Illustration of OOD recommendation. The distribution shift of user and item features in test phase, leading to decline in recommendation performance.
The out-of-distribution (OOD) problem has been extensively studied in various fields, including image classification [13] and detection [14]. However, it has received limited attention in the context of recommendation systems. To address this issue, disentangled recommendation approaches [15, 16] aim to learn factorized representations of user preferences, which can enhance the model robustness against distribution shifts. Causal-based methods [11, 12] use causal learning to solve OOD problems. However, these methods often require interventions during the training process, which makes them less practical when only a pre-trained model is available. Another approach to address OOD recommendation is model retraining [8], which involves adapting the model to the OOD environment. However, this approach often requires a substantial computational overhead. [11]. In contrast, our study aims to investigate practical approaches that can readily improve the generalization and robustness for a wide range of pre-trained models in different scenarios. Test-time training (TTT), a widely adopted strategy for solving OOD problems in various domains [17], serves as the inspiration for our proposed solution. The core concept of test-time training is to utilize auxiliary unsupervised or self-supervised learning (SSL) tasks to provide insights into the underlying characteristics of a test sample. This information is then used to fine-tune or calibrate the trained model specifically for that test sample. Since each sample has distinct properties, test-time training has significant potential to enhance the generalization capabilities of trained models.
To address the OOD recommendation challenge, we present a novel dual test-time-training strategy called DT3OR. This strategy leverages test-time training to adapt the model during the recommendation phase and improve its performance in OOD scenarios. In DT3OR, we introduce a model adaptation mechanism during the test-time phase to effectively update the recommendation model, thus allowing the model to adapt to the shift data. This mechanism consists of two key components to better understand the invariant preferences among users and the variant user/item features with shift data: the self-distillation task and the contrastive task. The self-distillation task aims to minimize the distance between user interest centers with the same preference, thereby improving the uniformity of user interest representations in the latent space. By encouraging the user interest centers to be more tightly clustered, the model enhances its proficiency in capturing the inherent patterns in user preference. Additionally, we incorporate a contrastive task that captures the correlations among users with similar preferences. This task involves selecting high-confidence samples from the same user interest center to establish meaningful relationships between users.
In summary, the main contributions of this work are summarized as follows:
  • To the best of our knowledge, this paper is the first work to deal OOD recommendation problem with test-time-training, which allows the model to adapt for the shifting user/item features.
  • To learn user-invariant preference relationships and enable the model to adapt to shifting user and item features, we propose a dual test-time training strategy that incorporates two self-supervised learning (SSL) tasks: the self-distillation task and the contrastive task.
  • We provide a theoretical analysis to demonstrate the rationale behind our designed dual test-time-training strategy.
  • Comprehensive experimental evidence across three datasets has shown that DT3OR surpasses other state-of-the-art recommendation algorithms in OOD scenario. In addition, the efficiency of the proposed modules is confirmed through ablation studies.
Figure 2: The overall framework for our proposed DT3OR.
Figure 3: Illustration of the self-distillation task designing. In our proposed method, we initially obtain the user embedding $\textbf{E}_{u}$ and item embedding $\textbf{E}_{i}$ by encoding the user/item features and historical interactions with the training network. After this, we carry out K-means on the combined embedding to yield the clustering results. Subsequently, with the sharpen function, we implement self-distillation to enhance the user interest centers, thus improving the generalization of model to deal with shift features.
II Problem Definition
OOD Recommendation Scenario. In this work, our primary focus is on addressing the out-of-distribution (OOD) problem in the recommendation task. We consider a training dataset denoted as $\mathcal{D}_{Tr}$, and a test dataset denoted as $\mathcal{D}_{Te}$. The corresponding set of interactions are denoted as $\mathcal{Y}_{Tr}$ and $\mathcal{Y}_{Te}$. Moreover, in an OOD scenario, $\mathcal{D}_{Tr}$ and $\mathcal{D}_{Te}$ are disjoint, and they are drawn from the different distribution.
Definition 1 (Test-time training for OOD Recommendation (T3OR).).
T3OR requires the learning of a self-supervised learning task denoted as $h(\cdot)$ during the test phase. By leveraging a pre-trained network $f_{\theta}$, T3OR aims to improve the performance on the test data $\mathcal{D}_{Te}$:
[…]
$s.t.\leavevmode\nobreak\ \leavevmode\nobreak\ h\left(\mathcal{D}_{Te}\right)\in\mathcal{P}\left(\mathcal{D}_{Te}\right)$
[…]
where $\mathcal{L}$ is the loss function to measure the performance of downstream task, i.e., recommendation. $\mathcal{P}\left(\mathcal{D}_{Te}\right)$ is the set of self-supervised tasks.
III Methodology
We design a novel framework termed DT3OR for recommendation under an OOD scenario in this section. The framework of DT3OR is shown in Fig. 2. To be specific, we first describe the overall framework of DT3OR. Then, we delve into the designed self-supervised learning tasks, which include the self-distillation task and the contrastive task. Finally, we give the definition of loss function in DT3OR.
The primary obstacle in optimizing the DT3OR problem lies in constructing a surrogate loss that can guide the learning process, particularly considering the lack of ground-truth interactions for the test data. To tackle this challenge, we present a dual self-supervised learning task as delineated below.
III-A Self-supervised Learning Tasks Designing
Our proposed DT3OR is designed to enhance model generalization and robustness to handle the OOD recommendations by learning from the test data. Ideally, if we had the ground-truth labels for the test data, we could refine the model to minimize the cross-entropy loss on the test data. However, due to the absence of such label information at the test-time, we are motivated to explore suitable surrogate loss functions to steer the model without the presence of labeled data. In the unsupervised scenario, self-supervised learning (SSL) techniques [18, 19, 20] have gained significant attention, laying the groundwork for self-supervision during test-time training. The effectiveness of test-time training heavily dependents on the creation of a relevant and informative SSL task [17]. However, SSL tasks commonly used in image-related tasks, such as rotation invariance, may not be directly applicable to recommendation data due to its distinct characteristics. In following, we design two SSL tasks that are suitable for recommendation tasks.
III-A1 Self-distillation Task Designing
In this subsection, we propose a test-time self-distillation task to make better leverage of the test data. To be specific, we first extract the embeddings of user/item features and historical interactions by the pretrain user feature network $f(\cdot)$ and item feature network $g(\cdot)$:
$\textbf{E}_{u}$ $=f(\textbf{X}_{u},\textbf{D})$ (2)
$\textbf{E}_{i}$ $=g(\textbf{X}_{i},\textbf{D}),$
where $\textbf{X}_{u}$ and $\textbf{X}_{i}$ denote the user features and item features, respectively. D is the historical interactions. $\textbf{E}_{u}$ and $\textbf{E}_{i}$ capture the preferences of users towards items. In a recommendation system, historical interactions and user features are valuable indicators of user preferences [21]. User’s purchase behaviors could reflect their preference for a particular type of item. Similarly, the items interacted with by users can indicate their preferences.
Then, we fuse the user embeddings $\textbf{E}_{u}$ and item embeddings $\textbf{E}_{i}$ by:
E $=\gamma(\textbf{E}_{u}\oplus\textbf{E}_{i}),$ (3)
where $\oplus$ denotes the the concatenation operation, $\gamma(\cdot)$ denotes a multi-layer perceptron (MLP). We implement K-means algorithm [22] on E. The clustering results distribution could be obtained by:
$q$ $=\varphi(\textbf{E}).$ (4)
where $\varphi(\cdot)$ denotes the K-means algorithm [22]. The variable $q$ can be interpreted as the probability of assigning a sample to a specific cluster, indicating a soft assignment. We can consider the matrix $\textbf{Q}=[q]$ as the distribution of assignments for all samples.
After obtaining the distribution of the clustering result Q, we utilize self-distillation to optimize the network by leveraging a lower entropy assignments. In particular, we employ a sharpening function to diminish the entropy of the clustering result distribution Q. We determine the target distribution P through the following calculation:
$\textbf{P}=S(q,T)_{i}$ $=\frac{q_{i}^{\frac{1}{T}}}{\sum_{j}q_{j}^{\frac{1}{T}}}.$ (5)
where $S(\cdot)$ represents the sharpening function, $T$ is a temperature parameter, and $q$ denotes input probability distribution, which corresponds to the clustering result distribution in this paper. As the temperature $T$ approaches zero, the output of $S(q,T)$ tends to approximate a Dirac distribution (one-hot encoding). A lower temperature encourages the generation of a lower-entropy distribution. Consequently, we can formulate the self-distillation objective function as follows:
$\mathcal{L}_{d}$ $=KL(\textbf{P}||\textbf{Q})=\sum_{i}p_{i}log\frac{p_{i}}{q_{i}}.$ (6)
By minimizing the KL divergence loss between the P and Q distributions, the target distribution P aids the model in acquiring better embeddings, given the low entropy of P. This process can be regarded as a self-distillation mechanism, considering that P is derived from Q. Furthermore, the P distribution guides the update of the Q distribution in a supervisory capacity.
Through clustering the fused embeddings of users and items, we obtain interest centers that represent the preferences of different types of users, where each type of user shows a preference for a specific type of item. The self-distillation mechanism designed in this paper enhances the prominence of these interest centers, allowing for better understanding and modeling of user preferences.
Figure 4: Illustration of the contrastive task designing. With the high confidence results of clustering, we select the samples in the same interest center as positive sample, while regarding the different high confidence interest centers as negative samples.
III-A2 Contrastive Task Designing
Contrastive learning has surfaced as a significant self-supervised framework, demonstrating promising performance across various tasks. In this subsection, we design a contrastive task specifically for test-time training. Data augmentation is pivotal in contrastive learning. However, inappropriate data augmentation can lead to semantic drift, thereby degrading the quality of positive samples [23]. Additionally, existing strategies for constructing negative samples often consider all non-positive samples as negative, which can inadvertently introduce false-negative samples [24].
In this subsection, we design a contrastive task with more discriminative positive samples and reliable negative samples. Based on the clustering results obtained by Eq.(4), we design a contrastive task in this section. To be concrete, the designed contrastive task contains two steps.
Firstly, we construct the positive sample pairs. Based on the clustering distribution $q$, we identify the high confidence sample indexes as:
$h=\text{Top}(q),$ (7)
where $\text{Top}(\cdot)$ is used to select the indices of the top $\tau$ high-confidence samples in the clustering distribution $q$. According to these indices, we denote these high-confidence embeddings as $\textbf{E}_{h}$. Subsequently, based on the pseudo labels derived from the clustering operation in Eq.(4), we partition $\textbf{E}_{h}$ into $K$ disjoint clusters $\textbf{R}_{i}$, which can be represented as:
$\textbf{R}_{i}=\text{Div}({\textbf{E}_{h}}),$ (8)
where $\text{Div}(\cdot)$ is the division function. $i\in[1,K]$. By this way, the high-confidence embeddings could be divided into different clusters, which denotes different preferences in user-item pairwise level. We consider the samples in the same clusters as positive samples. In other words, the same preference relationship (user-like-item) is considered a positive samples.
After that, we calculate the centers of high confidence embeddings to construct negative samples:
$\textbf{C}_{i}=\frac{1}{K}\sum_{i=1}^{K}{\textbf{R}_{i}},$ (9)
where $\textbf{C}_{i}$ represents different preference centers. We consider different high-confidence centers as negative samples, as they correspond to different interest centers. In this way, we could reduce the possibility of false-negative samples.
In conclusion, utilizing high-confidence clustering information as a form of supervised information could boost the quality of both positive and negative samples, thereby resulting in enhanced performance in contrastive tasks.
Algorithm 1 Dual Test-Time Training for OOD Recommendation
III-B Objective function
We define the loss function of our DT3OR in OOD recommendation in this subsection. DT3OR jointly optimizes two components, i.e., the self-distillation loss $\mathcal{L}_{d}$ and contrastive loss $\mathcal{L}_{c}$.
In detail, we utilize the Mean Squared Error loss between positive samples as the positive loss:
$\mathcal{L}_{p}=\frac{1}{KN(N-1)}{\sum_{m=1}^{K}\sum_{i=1}^{N}\sum_{j=i+1}^{N}{(\textbf{R}^{m}_{i}-\textbf{R}^{m}_{j})^{2}}},$ (10)
where $K$ is utilized to denote the number of interest clusters. $\textbf{R}^{m}_{i}$ and $\textbf{R}^{m}_{j}$ denotes the embeddings of different sample in the same cluster $\textbf{R}^{m}$. Besides, we use $N$ to present the number of high-confidence samples.
Moreover, we implement the negative sample loss with cosine similarity among different centers of the selected high-confidence embeddings. The initial step involves computing the cosine similarity between different centers:
$sim(\textbf{C}_{i},\textbf{C}_{j})=\frac{\textbf{C}_{i}\cdot\textbf{C}_{j}}{||\textbf{C}_{i}||_{2}||\textbf{C}_{j}||_{2}}.$ (11)
Then, we maximize the cosine similarity for different centers as:
$\mathcal{L}_{n}=\frac{1}{K^{2}-K}\sum_{i=1}^{K}\sum_{j=i+1}^{K}{{sim(\textbf{C}_{i},\textbf{C}_{j})}},$ (12)
where $\textbf{C}_{i}$ and $\textbf{C}_{j}$ denote different high confidence center. In summary, the contrastive loss $\mathcal{L}_{c}$ is:
$\mathcal{L}_{c}=\mathcal{L}_{p}+\mathcal{L}_{n}.$ (13)
The whole loss of DT3OR can be presented as follows:
$\mathcal{L}=\mathcal{L}_{d}+\alpha\mathcal{L}_{c},$ (14)
where we use $\alpha$ to represent the trade-off hyper-parameter between $\mathcal{L}_{d}$ and $\mathcal{L}_{c}$.
The first term in Eq.(14) promotes the distribution of user interest centers closer together, enhancing the uniformity of user interest representations in the latent space. The second term captures the correlations among users with similar preferences. Here, we use ${B}$ to represent the batch size. Furthermore, the embeddings’ dimension is set to $d$, and $N$ as the number of high-confidence samples. Consequently, we could obtain time complexity of the self-distillation loss and contrastive loss is $\mathcal{O}(Bd)$ and $\mathcal{O}(N^{2}d)$, respectively. Therefore, the total time complexity of the loss is $\mathcal{O}(Bd+N^{2}d)$. The space complexity of our loss is $\mathcal{O}(B+N^{2})$. The comprehensive learning process of DT3OR is outlined in Algorithm 1.
IV Theoretical analysis
[…]
$\mathcal{L}_{CE}=-\frac{1}{N}\sum_{i=1}^{N}y_{i}\text{log}(\hat{y}_{i}),$ (15)
where $y_{i}$ denotes the ground-truth. $\hat{y}_{i}$ is the prediction of the model embedding, which can be presented as:
$\hat{y}_{i}$ $=softmax(f(u,i))$ (16)
$=softmax(g_{i})$
$=\frac{e^{g_{i}}}{\sum_{j=1}^{N}e^{g_{j}}}$
where $f(\cdot)$ is the recommendation model, $y_{i}$ is defined as the variable indicating whether the user has interacted with an item, while $\hat{y}_{i}$ represents the predicted probability of interaction between the user and the item by the model.
Theorem 1.
$\mathcal{L}(u,i,y;\theta)$ is convex and $\beta$-smooth with respect to ${\theta}$, and the magnitude of the gradient $|\nabla{\bm{\theta}}\mathcal{L}(u,i,y;\theta)|$ is bounded by a positive constant $B^{{}^{\prime}}$ for all ${\bm{\theta}}$.
Proof.
With the inner product of embeddings $g=f(u,i)$ and based on Eq (15), we could have the following formula:
$\mathcal{L}(y;g)$ $=-\frac{1}{N}\sum_{i=1}^{N}(y_{i}\text{log}(\hat{y}_{i})$ (17)
$=-\frac{1}{N}\sum_{i=1}^{N}(y_{i}\text{log}(\frac{e^{g_{i}}}{\sum_{j=1}^{N}e^{g_{j}}}))$
where $\hat{y}_{i}$ is the prediction of the recommendation model. Then, we calculate the Hessian Matrix of Eq (17):
$if\leavevmode\nobreak\ \text{i}\neq\text{j},$ $\leavevmode\nobreak\ \mathcal{\bf M}_{ij}=-\hat{y}_{i}\hat{y}_{j},$
$if\leavevmode\nobreak\ \text{i}=\text{j},$ $\leavevmode\nobreak\ \mathcal{\bf M}_{ij}=\hat{y}_{i}-\hat{y}_{i}^{2}.$
Based on Eq.(IV), for any ${\bf c}\in\mathbb{R}^{N}$, we could have:
Based on Cauchy-Schwarz Inequality [25], we scale the above equation:
$(\sum_{i=1}^{N}{\bf c}_{i}\sqrt{\hat{y}_{i}}\sqrt{\hat{y}_{i}})^{2}\leq(\sum_{i=1}^{N}{{\bf c}_{i}}^{2}\hat{y}_{i})(\sum_{i=1}^{N}\hat{y}_{i}).$
Therefore,
${\bf c}^{\mathsf{T}}\mathcal{\bf M}{\bf c}$ $\geq\sum_{i=1}^{N}\hat{y}_{i}{\bf c}_{i}^{2}-(\sum_{i=1}^{N}{{\bf c}_{i}}^{2}\hat{y}_{i})(\sum_{i=1}^{N}\hat{y}_{i}).$
Since $\hat{y}_{i}$ is obtained by the softmax, $\sum_{i=1}^{N}\hat{y}_{i}=1.$. Therefore:
${\bf c}^{\mathsf{T}}\mathcal{\bf M}{\bf c}$ $\geq 0.$
Therefore, we have successfully demonstrated that the Hessian Matrix $\mathcal{\bf M}$ of Eq (17) is positive semi-definite (PSD). This proof is equivalent to establishing the convexity of $\mathcal{L}(y;{\bf g})$ with respect to $\bf g$.
For any ${\bf r}\in\mathbb{R}^{N}$ and $\|{\bf r}\|=1$, we could obtain:
${\bf r}^{\mathsf{T}}\mathcal{\bf M}{\bf r}$ $=\sum_{i=1}^{N}\hat{y}_{i}{\bf r}_{i}^{2}-(\sum_{i=1}^{N}{{\bf r}_{i}}\hat{y}_{i})^{2}$
$=\sum_{i=1}^{N}\hat{y}_{i}{\bf r}_{i}^{2}-(\sum_{i=1}^{N}{\bf r}_{i}\sqrt{\hat{y}_{i}}\sqrt{\hat{y}_{i}})^{2}$
$\leq\sum_{i=1}^{N}{\bf r}_{i}^{2}+(\sum_{i=1}^{N}{\bf r}_{i}\sqrt{\hat{y}_{i}}\sqrt{\hat{y}_{i}})^{2}.$
Similarly, with Cauchy-Schwarz Inequality [25], we have:
${\bf r}^{\mathsf{T}}\mathcal{\bf M}{\bf r}$ $\leq\sum_{i=1}^{N}{\bf r}_{i}^{2}+(\sum_{i=1}^{N}{\bf r}_{i}\sqrt{\hat{y}_{i}}\sqrt{\hat{y}_{i}})^{2}$
$\leq\sum_{i=1}^{N}{\bf r}_{i}^{2}+(\sum_{i=1}^{N}{\bf r}_{i}^{2})(\sum_{i=1}^{N}\hat{y}_{i}).$
Since $\|{\bf r}\|=1$ and $\sum_{i=1}^{N}\hat{y}_{i}=1$, we have
${\bf r}^{\mathsf{T}}\mathcal{\bf M}{\bf r}$ $\leq 2.$
Therefore, we have successfully proved that the eigenvalues of the Hessian Matrix $\mathcal{\bf M}$ are bounded by 2. This proof is equivalent to establishing the $\beta$-smoothness of $\mathcal{L}(y;{\bf g})$ with respect to $\bf g$.
After that, we calculate the first-order gradient of of $\mathcal{L}(y;{\bf g})$ over $\bf{g}$:
$if\leavevmode\nobreak\ \text{i=c},\nabla_{\bf g}{\mathcal{L}(u,i,{y};{\bf g}})=-1+\hat{y}_{i},$ (18)
$if\leavevmode\nobreak\ \text{i}\neq\text{c},\nabla_{\bf g}{\mathcal{L}(u,i,{y};{\bf g}})=\hat{y}_{i}.$
We implement $\ell_{2}$ on Eq.(18):
$\|\nabla_{\bf g}{\mathcal{L}({y};{\bf g}})\|$ $=\sqrt{(-1+\hat{y}_{c})^{2}+\sum_{i\neq c}(\hat{y}_{i})^{2}}$
$=\sqrt{\sum_{i\neq c}(\hat{y}_{i})^{2}+\sum_{i\neq c}(\hat{y}_{i})^{2}}$
$\leq 2.$
Consequently, we have established that the $\ell_{2}$ norm of the first-order gradient of $\mathcal{L}(y;{\bf g})$ is bounded by a positive constant.
By demonstrating the convexity and $\beta$-smoothness of $\mathcal{L}(y;{\bf g})$ for all ${\bf g}$ and $y$, alongside the condition $|\nabla_{\bf g}\mathcal{L}(y;{\bf g})|\leq B$ for all ${\bf g}$, where $B$ is a positive constant, we have successfully completed the proof of Theorem 1.
Theorem 2.
Assume that for all $(u,i,y)$, $\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})$ is differentiable, convex and $\beta$-smooth in ${\bm{\theta}}$, and both
$\|\nabla_{{\bm{\theta}}}\mathcal{L}_{SSL}(u,i;{\bm{\theta}})\|\leq B$ for all ${\bm{\theta}}$, where $B$ is a positive constant. With a fixed learning rate $\delta=\frac{\epsilon}{\beta B^{2}}$, for every $(u,i,y)$ such that
$\left\langle\nabla_{{\bm{\theta}}}\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}),\nabla_{{\bm{\theta}}}\mathcal{L}_{SSL}(u,i;{\bm{\theta}})\right\rangle>\epsilon,$
Thus,
$\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})>\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}(u,i)),$
where ${\bm{\theta}}(u,i)={\bm{\theta}}-\delta\nabla_{{\bm{\theta}}}\mathcal{L}_{SSL}(u,i;{\bm{\theta}})$.
Remark. By applying Theorem1, we can assert that the objective loss function of the recommendation model possesses the properties of differentiability, convexity, and $\beta$-smoothness. Furthermore, in combination with TheoremIV, we can conclude that test-time training could enhance the performance of the recommender system.
Proof.
Following [26], we prove Theorem 2 in this section. Specifically, for any $\delta$, by $\beta$-smoothness, we could obtain:
$\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}(u,i)})$ $=\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}(x)}-\delta\nabla_{\bm{\theta}}{\mathcal{L}_{SSL}(u,i;{\bm{\theta}})})$
$\leq\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}(u,i)}$
$-\delta\left\langle\nabla_{{\bm{\theta}}}\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}),\nabla_{{\bm{\theta}}}\mathcal{L}_{SSL}(u,i;{\bm{\theta}})\right\rangle$
$+\frac{\delta^{2}\beta}{2}\|\nabla_{\bm{\theta}}{\mathcal{L}_{SSL}(u,i;{\bm{\theta}})}\|^{2}.$
Then, we denote $\delta^{*}$ as follows:
$\delta^{*}=\frac{\left\langle\nabla_{{\bm{\theta}}}\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}),\nabla_{{\bm{\theta}}}\mathcal{L}_{SSL}(u,i;{\bm{\theta}})\right\rangle}{\beta\|\nabla_{\bm{\theta}}{\mathcal{L}_{SSL}(u,i;{\bm{\theta}})}\|^{2}},$
Thus, we could calculate as:
$\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}-\delta^{*}\nabla_{\bm{\theta}}{\mathcal{L}_{SSL}(u,i;{\bm{\theta}})})$
$\leq\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})-\frac{{\left\langle\nabla_{{\bm{\theta}}}\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}),\nabla_{{\bm{\theta}}}\mathcal{L}_{SSL}(u,i;{\bm{\theta}})\right\rangle}^{2}}{2\beta\|\nabla_{\bm{\theta}}{\mathcal{L}_{SSL}(u,i;{\bm{\theta}})}\|^{2})}.$
Based on the given assumptions regarding the $\ell_{2}$ norm of the first-order gradients for the main task and the self-supervised learning (SSL) task, as well as the assumption on their inner product, we can derive the following:
$\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})$ $-\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}-\delta^{*}\nabla_{\bm{\theta}}{\mathcal{L}_{SSL}(u,i;{\bm{\theta}})})$
$\geq\frac{\epsilon^{2}}{2\beta B^{2}}.$
Here, we denote $\delta=\frac{\epsilon}{\beta B^{2}}$. Besides, based on the assumptions. we could obtain $0<\delta\leq\delta^{*}$.
Then, denote ${\bf t}=\nabla_{\bm{\theta}}{\mathcal{L}_{SSL}(u,i;{\bm{\theta}})},$ by convexity of $\mathcal{L}_{SUP}$,
$\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}(u,i))=\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}-\delta{\bf t})$
$=\mathcal{L}_{SUP}(u,i,y;(1-\frac{\delta}{\delta^{*}}){\bm{\theta}}+\frac{\delta}{\delta^{*}}({\bm{\theta}}-\delta^{*}{\bf t}))$
$\leq(1-\frac{\delta}{\delta^{*}})\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})+\frac{\delta}{\delta^{*}}\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}-\delta^{*}{\bf t})$
$\leq(1-\frac{\delta}{\delta^{*}})\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})+\frac{\delta}{\delta^{*}}(\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})-\frac{\epsilon^{2}}{2\beta B^{2}})$
$=\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})-\frac{\delta}{\delta^{*}}\frac{\epsilon^{2}}{2\beta B^{2}}.$
Moreover, $\frac{\delta}{\delta^{*}}>0$, we could calculate:
$\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}})-\mathcal{L}_{SUP}(u,i,y;{\bm{\theta}}(u,i))>0.$
Therefore, we can conclude that test-time training could enhance the performance of the recommender system.
Table I: Statistics of the datasets used in our method. Besides, ”int” represents interactions.
Dataset Synthetic Meituan Yelp
#User 1,000 2,145 7,975
#Item 1,000 7,189 74,722
#OOD Int 112,371 6,944 99,525
Density 0.257641 0.001189 0.000679
V Experiments
We implement a series of experiments in this section to validate the effectiveness of the proposed DT3OR by addressing the following research questions:
  • RQ1: How does the performance of DT3OR in OOD recommendation compared with other SOTA recommender methods?
  • RQ2: What is the impact of the hyper-parameters on the performance of DT3OR?
  • RQ3: How do the distinct components of DT3OR contribute to its performance?
  • RQ4: What is the underlying interest clustering structure revealed by DT3OR?
Table II: OOD recommendation performance comparison on three datasets with seven baselines. Especially, the top two results are emphasized in bold and underlined, denoting the best and second-best results, respectively. Besides, $\dagger$ denotes results are statistically significant where the p-value is less than 0.5.
Dataset Synthetic Data Meituan Yelp
Metricc R@10 R@20 N@10 N@20 R@50 R@100 N@50 N@100 R@50 R@100 N@50 N@100
FM 5.72 10.74 6.04 7.92 1.21 2.05 0.43 0.57 9.64 13.89 3.13 3.85
NFM 4.05 7.61 4.38 5.60 2.33 3.54 0.66 0.85 8.29 12.76 2.41 3.16
MultiVAE 2.08 4.08 1.72 2.57 2.38 3.68 0.69 0.91 3.65 5.82 1.18 1.54
MacridVAE 2.31 3.92 1.92 2.62 2.19 3.64 0.67 0.90 4.08 6.34 1.35 1.74
MacridVAE+FM 4.63 8.36 5.13 6.43 2.33 3.64 0.66 0.87 4.07 6.26 1.40 1.78
CausPref 5.41 4.77 6.74 9.41 3.44 4.35 0.98 1.21 13.46 18.34 4.14 5.01
COR 7.67 14.43 8.04 10.56 3.68 5.78 1.01 1.35 14.16 19.86 5.00 5.95
Ours 8.20 14.84 8.42 10.86 4.15 5.83 1.13 1.40 14.41 19.90 5.10 6.02
%Improvement 6.91% 2.84% 4.73% 2.84% 12.77% 0.87% 11.88% 3.70% 1.77% 0.20% 2.00% 1.18%
Synthetic
Meituan
Yelp
Figure 5: Sensitivity analysis of the hyper-parameter cluster number $K$ on Synthetic, Meituan and Yelp datasets.
V-A Datasets & Metric
V-A1 Benchmark Datasets.
Following COR [11], we conduct extensive experiments to evaluate the effectiveness of DT3OR on three benchmark datasets, including Meituan  [^1] https://www.biendata.xyz/competition/smp2021_2/, Yelp  [^2] https://www.yelp.com/dataset., and one synthetic dataset.
1) The synthetic dataset contains 1,000 users and 1,000 items. The features of the users/items are sampled from the standard Gaussian $\mathcal{N}(0,1)$. The OOD features are collected by re-sampling user/item features from $\mathcal{N}(1,1)$. The detailed description of the dataset construction is provided in Appendix. LABEL:Constrcution.
2) Meituan includes a public food recommendation dataset that contains a large number of user and item features, e.g., user consumption levels and food prices. For the OOD scenario in Meituan, we define it as a shift in the average consumption levels between weekdays and weekends. Specifically, users exhibit higher or lower consumption levels during weekends compared to weekdays. The user interactions during weekdays are considered in-distribution interactions, while those during weekends are considered OOD interactions.
3) Yelp is a commonly utilized dataset for restaurant recommendations. Within this dataset, the user’s location is taken into account as the fluctuating feature. The user ratings are sorted by timestamps, and the dataset is divided into two parts based on changes in location. The user interactions with shifted locations are considered OOD interactions.
For all datasets, interactions that have ratings of 4 or higher are treated as positive instances. To maintain a level playing field, a sparse partitioning approach is employed, allocating 80%/10%/10% of the data for training, validation, and testing. Comprehensive statistics of the datasets used in this paper are listed in Table I.
V-A2 Evaluation Metrics.
The performance of the recommendation is evaluated through two commonly applied metrics, i.e., Recall and NDCG. These metrics are evaluated using the all-ranking protocol [27], which considers the top-K items selected from the complete set of items that users have not interacted with. For the synthetic dataset, we designate K values of 10 and 20. Given the significant quantity of items in the Meituan and Yelp datasets, we showcase the results for K values of 50 and 100.
V-B Experiment Settings
V-B1 Implementation Details.
The experiments are executed utilizing an NVIDIA 3090 GPU on the PyTorch platform. For the benchmark models, the performance values from the COR [11] paper are directly reported for all datasets. In our proposed approach, we utilized COR as the primary network and trained our network using the Adam optimizer [28]. The batch size was established at 500. Moreover, the learning rate and dropout rate were set at 1e-4 and 0.5 respectively for all datasets. The training during test-time was carried out for 10 epochs. The trade-off hyper-parameter $\alpha$ was selected from the set $\{0.01,0.1,1.0,10,100\}$. The number of clusters $K$ was consistently set at 4, and the threshold for selecting high confidence samples was maintained at 0.9.
V-B2 Comparison Recommendation Methods.
To substantiate the efficiency of our DT3OR, we contrast its performance with several baseline models, including FM [29], NFM [30], MultiVAE [31], MacridVAE [32], MacridVAE+FM, CausPref [12], and COR [11]. Specifically,
  • FM [29]: With factorized parameters, FM models all interactions between variables.
  • NFM [30]: NFM fuses the linearity of FM in handling second-order feature interactions with the non-linearity exhibited by networks.
  • MultiVAE [31]: MultiVAE used variational autoencoders to collaborative filtering for the implicit feedback.
  • MacidVAE [32]: It acquires disentangled representations by deducing high-level concepts linked to user intentions.
  • CausPref [12]: CausPref used causal learning to obtain the invariant user preference and anti-preference negative sampling to handle the implicit feedback.
  • COR [11]: COR devised a new variational autoencoder to deduce the unobserved features from historical interactions in a causal way with out-of-distribution recommendation.
V-C (RQ1) Performance Comparisons
We evaluate the experiments against seven baseline models based on Recall and NDCG metrics. These models fall into three classes, which include FM-based methods (FM [29], NFM [30]), VAE-based methods (MultiVAE [31], MacridVAE [32], MacridVAE+FM), and causal-based methods (CausPref [12], COR [11]).
The comparison results are demonstrated in Table. II. Observations drawn from the results are as follows:
1) Our proposed DT3OR could achieve SOTA recommendation performance compared to classical FM-based recommendation algorithms. We infer that the reason behind this is that these methods do not design specific strategy to handle the user/item feature shift;
2) FM-based recommendation algorithms have a significant advantage on datasets with notable feature shifts, resulting in better performance compared to VAE-based algorithms on synthetic and Yelp datasets;
3) The test-time training with SSL tasks in our method leads to improved performance compared to causal-based recommendation methods.
In summary, our proposed DT3OR outperforms several recommendation methods across various datasets in the OOD scenario. Taking the results on meituan dataset for example, DT3OR exceeds the second-best method by margins of 12.77%, 0.87%, 11.88%, and 3.70% in R@50, R@100, N@50, and N@100, respectively. The test-time-training with dual SSL tasks could assist the model adapt to the shift user/item features in test-time phase under OOD environment, thus alleviating the influence of data shift.
V-D (RQ2) Parameter Sensitivity
We implement series of parameter analyses on Meituan, Yelp and the Synthetic datasets to verify the sensitivity of the hyper-parameters in this subsection.
Synthetic
Meituan
Yelp
Figure 6: Sensitivity analysis for the trade-off hyper-parameter $\alpha$.
Synthetic
Meituan
Yelp
Figure 7: Sensitivity analysis for the threshold hyper-parameter $\tau$.
Synthetic with R@10
Meituan with R@50
Yelp with R@50
Synthetic with R@20
Meituan with R@100
Yelp with R@100
Synthetic with N@10
Meituan with N@50
Yelp with N@50
Synthetic with N@20
Meituan with N@100
Yelp with N@100
Figure 8: Ablation studies over proposed DT3OR on three datasets. “(w/o) SSL1”, “(w/o) SSL2” and “(w/o) SSL1&SSL2” denote the reduced models by individually removing the self-distillation task, the contrastive task, and all aforementioned modules combined.
V-D1 Sensitivity Analysis for trade-off $\alpha$
To investigate the impact of the trade-off parameters $\alpha$ in our model, we performed experiments on the Synthetic, Meituan, and Yelp datasets. We varied the values of $\alpha$ within the range $\{0.01,0.1,1.0,10,100\}$. It is to be noted that all experimental settings remained consistent across different values of $\alpha$. The results of these experiments are depicted in Fig. 6. The observations are as follows.
1) When $\alpha$ is assigned extreme values, such as $0.01$ or $100$, the performance of the recommendation system tends to decrease.
2) Extreme values can disrupt the balance between different loss components, whereas values around $1.0$ tend to yield better results.
K = 3
K = 4
K = 5
K = 6
Figure 9: Visualization experiment on Synthetic dataset with different cluster number K.
V-D2 Sensitivity Analysis for cluster number $K$
We implement sensitive experiments to explore the impact of the parameter $K$, which represents the number of interest centers. We varied the value of $K$ within the range of $\{2,4,10,20,100\}$. According to the results in Fig. 5, we could obtain the following observations.
1) The model achieves the best recommendation performance when $K$ is around 4. When $K$ takes extreme values, e.g., $K=100$, the performance will decrease dramatically. We speculate that this is because the interest centers become too scattered, making it difficult to accurately reflect the true preferences of users.
2) A similar situation occurs when $K=2$, where having too few interest centers fails to effectively capture the diverse preferences of users.
V-D3 Sensitivity Analysis for the threshold $\tau$
We explore the impact of the threshold $\tau$ by implementing experiments on three datasets. The results are presented in Fig.7. We could conclude that the model achieves promising performance as $\tau$ increases. This can be attributed to the fact that the high-confidence centers $\textbf{C}_{i}$ become more reliable with a higher threshold.
V-E (RQ3) Ablation Studies
We implement ablation studies in OOD recommendation scenario to validate the effectiveness of the SSL tasks we designed: self-distillation task and contrastive task. Concretely, “(w/o) SSL1”, “(w/o) SSL2” and “(w/o) SSL1&SSL2” denote the reduced models where we individually removed the self-distillation task, the contrastive task, and all aforementioned modules combined. Here, we adopt COR [11] as the backbone network to acquire representations for recommendation task, referred to as “(w/o) Cau&Con”. We have only removed the ablation module section while keeping the other modules unchanged from the DT3OR for conducting the ablation experiments. Specifically, apart from removing the corresponding modules, we keep identical parameter settings and training processes as used in the DT3OR.
The results illustrated in Fig 8 indicate that the removal of any of the designed modules results in a significant decrease in recommendation performance, which indicate the contribution of each module to the model’s overall efficacy. We delve deeper into the reasons for these results as follows:
1) We utilize clustering task to obtain the interest centers of different types of users, which could better benefit the model to learn the preferences (user-like-item) in test-time phase of shift user/item features.
2) Benefiting from the data mining capacity, the contrastive mechanism could improve the discriminative capacity of the model, thus reducing the impact of the shift in user/item features.
V-F (RQ4) Visualization Analysis
To reveal the intrinsic interest clustering structure, we utilize visualization experiments to display the distribution of the learned embeddings in this subsection. To be concrete, experiments are implemented with $t$-SNE algorithm [33] on the Synthetic dataset. We perform visualization experiments with varying cluster numbers, i.e., $K\in[3,4,5,6]$. As presented in Figure 9, the visual results emphasize that our proposed DT3OR method effectively identifies and represents the underlying interest clusters.
VI Related Work
VI-A Test-time Training
Test-time training (TTT)[17, 34, 35] is a method to improve model generalization in recent years, which allows for partial adaptation of a model based on test samples. TTT addresses the issue of distribution shifts between the training and test sets by updating the model using self-supervised learning (SSL) tasks during the test-time phase. Initially, TTT was introduced to update the model by jointly optimizing the supervised loss and SSL loss during the test period. However, these paradigms necessitate the simultaneous optimization of the supervised loss and SSL loss. To surmount this constraint, Tent [36] presents a fully test-time training approach that solely uses test samples and a pre-trained model. Inspired by the success of TTT in handling out-of-distribution problems, we apply this technique to the OOD recommendation scenario. By leveraging TTT, we can adapt the recommendation model using test samples and a pre-trained model, improving its performance in the presence of distribution shifts.
VI-B Self-Supervised Learning
Self-supervised learning (SSL) algorithms have recently garnered significant attention in the domain of graph neural networks (GNNs). SSL is a widely employed technique for training neural networks as evidenced in various studies [37, 38, 19, 39, 40]. A fundamental technique in SSL is contrastive learning, which focuses on increasing the similarity of positive pairs while decreasing the similarity of negative pairs. This approach has achieved remarkable success in graph learning [41, 42, 43, 44]. For instance, DCRN [45] strengthens feature decorrelation by aligning the representations of identical nodes across varied views while distancing the representations of dissimilar nodes in different views. Similarly, GMI [46] and HDMI [47] extend DGI [48] by incorporating edges and node features, to mitigate representation collapse. Additionally, the efficacy of amplifying mutual information to extract and learn embeddings for graph classification tasks has been demonstrated by MVGRL [49] and InfoGraph [50]. Building on these principles, GraphCL [51] and GRACE [52] produce two augmented views for SSL, and learn the representations by aligning the same nodes in both views and separating distinct nodes.
VI-C Recommender System
Recommender systems [53, 54, 35] have become a vital tool for discovering user’s interests and preferences in various online platforms, which assume that the user and item follow the independent and identically distributed (I.I.D.). OOD recommendation has received limited attention, which can be roughly divided into three classes. 1) Disentangled recommendation approaches [15, 16] aim to learn factorized representations of user preferences to enhance robustness against distribution shifts. 2) Causal-based methods [11, 12] leverage causal learning techniques to address OOD problems. However, these methods often require interventions during the training process, which may be impractical when only a pre-trained model is available. 3) Model retraining [8] focuses on adapting the model to the OOD environment. This approach necessitates a trade-off between retraining frequency and computational cost. Benefiting from the success of test-time training (TTT), we introduce the TTT strategy to assist model adaptation in OOD scenarios in this paper.
VII Conclusion
In this paper, we first introduce the test-time training strategy to address the out-of-distribution recommendation problem. The model could adapt the shift user/item feature distribution in test-time phase with our carefully designed self-supervised learning tasks. Specifically, the proposed self-distillation task and the contrastive task assist the model learn user’s invariant interest performance and the shift feature distributions in the test-time phase, resulting in the model achieving better performance on test data. Moreover, we give theoretical analysis to support the rationale of our dual test-time training strategy. Comprehensive experiments have affirmed the efficacy of DT3OR.